On January 1, 2021, California Financial purchases a building for $900,000, signing a 5%, 20-year mortgage. Installment payme1. Record the issuance of a note payable on Jan 01, 2021.

2. Record the first monthly mortgage payment on January 31, 2021.

3. Record the second monthly mortgage payment on February 28, 2021.

4. If total payments over the 20 years are $1,425,504 ($5,939.60 × 240 monthly payments), how much of this is interest expense and how much is actual payment of the loan? Actual payment = ? Interest expense = ?
Answer #1
1
January 01,2021 Building 900000
       Notes payable 900000
January 31,2021 Interest expense 3750.00 =900000*5%/12
Notes payable 2189.60
       Cash 5939.60
February 28,2021 Interest expense 3740.88 =(900000-2189.6)*5%/12
Notes payable 2198.72
       Cash 5939.60
4
Actual payment 900000
Interest expense 525504 =1425504-900000
Note: Alternatively Mortgage payable account may be used instead of Notes Payable